OpenStudy (firejay5):

Describe the change to the parent function. I am not good at this and I don't really know how to describe each parent function. 1. y = -3sqrt x 2. y = 2/5sqrt x 3. y = sqrt -x + 3 4. y = sqrt x - 2 5. y = sqrt x + 2 6. y = sqrt x - 4 7. y = 2 sqrt -x + 1 8. y = - sqrt x - 5 9. y = 4 sqrt x + 3 -1 10. y = -1/3 sqrt x -1 + 3

OpenStudy (jdoe0001):

hmm doyou have a parent function to start with?

OpenStudy (firejay5):

the parent function is obvious: y = sqrt x

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

so let's use that \(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards\\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ % template start \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ A}} \sqrt{x+ {\color{red}{ C}}} + {\color{green}{ D}}\\ %template end \qquad\qquad \quad \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ {\color{red}{ C}}\end{array} \begin{array}{llll}{\color{red}{ C}} > 0 & to\ the\ left\\ {\color{red}{ C}} < 0& to\ the\ right\end{array} \end{array} }\) see the shifts and expansions or shrinking?

OpenStudy (jdoe0001):

a negative leading "A" value, means, the function is flipped upside down

OpenStudy (firejay5):

don't forget B

OpenStudy (jdoe0001):

hmmm I do have a "B", in a more general template it's just no needed in this case

OpenStudy (jdoe0001):

so let's see the 1st 3 \(\bf y={\color{purple}{ -3}}\sqrt{x}\qquad y={\color{purple}{ \cfrac{2}{3}}}\sqrt{x}\qquad \sqrt{-x{\color{red}{ +3}}}\) actually the 3rd one, uses B so lemme post that for that one then

OpenStudy (jdoe0001):

\(\bf y={\color{purple}{ -3}}\sqrt{x}\qquad y={\color{purple}{ \cfrac{2}{3}}}\sqrt{x}\qquad \sqrt{{\color{blue}{ -1}}x{\color{red}{ +3}}}\) and check the template \(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards \\ \quad \\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ % template start y = {\color{purple}{ A}} ( {\color{blue}{ B}}x + {\color{red}{ C}} ) + {\color{green}{ D}}\\ % template ends \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to\ the\ left \\ \quad \\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to\ the\ right\end{array} \end{array} }\)

OpenStudy (jdoe0001):

anyhow,, need to dash... but you can see the 1st one is flipped upside down, and shrinks by a factor of 3 2nd one expands by a factor of 2/3 the bigger the number, it shrinks the graph, the smaller, it expands it the 3rd one has a horizontal shift of 3/-1 or -3 A* B means 1 * -1 => -1, so it's also flipped upside down

OpenStudy (firejay5):

@satellite73 shouldn't #1 be: refected over x-axis and expansion by a factor of 3